OpenStudy (anonymous):

(x^2+4x-8) divide by (2x) I got the answer x^2+3x-8 please advise?

OpenStudy (dumbcow):

\[\frac{x^{2}+4x-8}{2x} = \frac{x^{2}}{2x} + \frac{4x}{2x}-\frac{8}{2x}=\frac{x}{2}+2-\frac{4}{x}\]

OpenStudy (anonymous):

same